A man is 24years older than his son.in e years times he will twice as old as his son now. How old is his father



1
Expert's answer
2021-01-16T17:22:02-0500

"e=2.718"


"(x+24)+2.718=2(x+2.718)\to x+26.718=2x+5.436"


"x=21.28(years)" . Answer